### v1

# Sales Whisperer: A Human-Inconspicuous Attack on LLM Brand Recommendations

Abstract
Large language model (LLM) users might rely on others (e.g., prompting services), to write prompts. However, the risks of trusting prompts written by others remain unstudied. In this paper, we assess the risk of using such prompts on brand recommendation tasks when shopping. First, we found that paraphrasing prompts can result in LLMs mentioning given brands with drastically different probabilities, including a pair of prompts where the probability changes by . Next, we developed an approach that can be used to perturb an original base prompt to increase the likelihood that an LLM mentions a given brand. We designed a human-inconspicuous algorithm that perturbs prompts, which empirically forces LLMs to mention strings related to a brand more often, by absolute improvements up to . Our results suggest that our perturbed prompts, 1) are inconspicuous to humans, 2) force LLMs to recommend a target brand more often, and 3) increase the perceived chances of picking targeted brands.
1 Introduction
With recent advances in LLMs, chatbots are becoming a ubiquitous part of users’ digital experience. Users interact and control chatbots through natural language (i.e., prompts), for a nearly endless number of tasks. However, despite the natural language interface, effective prompts are often hard to create, leading some researchers to develop prompt optimization techniques (e.g., [1]). The industry has adopted the same ideas and has developed methods to recommend useful prompts to users. Chatbot services may suggest prompts for users (see Fig. 1), and dedicated forums may suggest prompts for users to try (see Fig. 2) [2]. While these services may convenient to users, little research has focused on implications of prompts created by other (untrusted) parties. Existing work has only explored various security aspects of LLMs [3, 4, 5, 6, 7, 8], but the risks of using prompts by others remain unexplored.
In this paper, we study whether inconspicuous manipulation of prompts can lead to LLM responses with substantial biases of an attacker’s choosing. We examine this problem by taking a first look at brand recommendation tasks: users may ask LLMs to recommend brands (e.g., Samsung) when shopping for a specific category of products (e.g., TVs). Adversaries, motivated financially, may aim to write and trick users into utilizing prompts that cause LLMs to recommend a specific brand more often. Similar to advertisements, adversaries gain economic benefits when a specific brand is recommended more often. In this scenario, such prompts and corresponding LLM response need to be human-inconspicuous to LLM users (i.e., users should not notice an attack is taking place) [10, 11, 12]: if users feel suspicious about either the prompts or the responses, they may not trust the prompt recommender and thus thwart the attack. We explain our threat model in detail in §3. While this study focuses on this specific attack scenario, the methods we describe in our work have the potential to be applied outside the realm of brand recommendation.
While many existing approaches have claimed to be capable of perturbing natural-language sequences (e.g., LLM prompts) in a human-inconspicuous manner [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 21, 24], these studies suffer one of two shortcomings: either their inconspicuousness isn’t evaluated via user studies or the approaches were human-detectable in accompanying user studies. Thus, in this work we take a different approach to perturbing prompts, ensuring that the perturbations are human-inconspicuous via an extensive user study. Further, our attack does not require access to the LLM weights or gradients, differentiating our work from most prior studies.
As a first step in evaluating potential attacks in our threat model, we measure how paraphrased prompts can result in drastically different LLM responses. We perform tests on Gemma-it (instruction-tuned), Llama2, Llama3, and Llama3-it (instruction-tuned). These models are described further in §5.1. To do so, we need a dataset of prompts for brand recommendation tasks. To the best of our knowledge, there were no public datasets for this purpose. Hence, we created a new dataset, consisting of 77 product categories and 449 prompts. In our experiments, we used the frequency of LLMs mentioning certain brands 111In practice, we target a set of words related to the brand (e.g., “Macbook,” “Apple” for the “Apple” brand). Throughout the paper, we refer to responses that contain any of these target words as the LLM mentioning said brand. as a proxy metric of how often LLMs recommend the brand. We later show that this proxy proved to be accurate (see 7). We observed that paraphrased alternatives of prompts result in a wide range of LLM responses, and can change the likelihood that LLMs mention a brand by up to in the most extreme case. On average, the difference between likelyhoods of two rephrased prompts ranged from and , depending on the LLM (§6.1).
The observation that similar prompts can result in different likelihoods of brand being mentioned motivates our next contribution. With access LLM to logits (notably, not weights), we propose a new approach that perturbs base prompts through synonym replacement. These human-inconspicuous perturbed prompts (§4.2) increase the likelihood of a target brand being mentioned in the LLM response. Candidate perturbed prompts are generated by replacing certain words with known synonyms, and one is selected based on a logit-based loss function. While several synonym dictionaries exist [25, 26, 23, 27, 28], we found that synonyms suggested by these dictionaries are human-detectable in the context of shopping brand recommendations, and thus we create our own synonym dictionary. We show that our method of performing synonym replacement can increase the likelihood that LLMs mention a brand by absolute improvements up to (§6.2). Across categories, the average best improvement between base prompt for that category and a perturbed prompt was , , , and for Gemma-it, Llama2, Llama3, and Llama3-it, respectively.
We also investigated the transferability [29, 30] of our synonym-replacement approach. We observed that the synonym replacement approach is highly transferable between a limited pair of open-sourced LLMs and GPT3.5, a close-sourced commercial LLM (§6.3).
Theoretical success in our attacks may not translate to real-world outcomes. To evaluate the effectiveness of our attacks in a more realistic setting, we conducted an extensive user study (§7). We specifically measure if participants will (1) find differences between perturbed/unperturbed prompt pairs, (2) find differences in responses to these pairs, and (3) be influenced by the increased likelihood of brand appearance in responses. We found that our synonym replacement attack acheives all three adversarial goals with statistical significance (§7.2.2), validating our earlier experiments.
In summary, our contributions are the following:
We collected a dataset of 449 prompts asking LLMs for recommendations on brands,
We identified a new threat model where adversaries perturb prompts and convince users to use them, ultimately causing LLMs to mention target brands.
We measured the difference in the likelihood of mentioning target brands in LLM’s responses to paraphrased prompts, finding a wide range.
We proposed a novel attack, synonym replacement approach to increase the likelihood of LLMs mentioning target brands. Through extensive evaluation, we show that this attack is successful.
We evaluated the transferability of this attack from open-souced LLMs to GPT, finding that it is transferable for a limited set of models.
Finally, through a user study, we showed that synonym replacement meets adversarial goals in a realistic setting.
The rest of the paper has the following layout: We give an overview of related work in §2. We define our threat model in §3. We then describe two approaches for finding prompts that can increase the probability of a brand being mentioned that adversaries may use in the new threat model, in §4 and the technical setups to evaluate these approaches in §5. We show our empirical results in §6. We describe our user study and its results in §7. Finally, we conclude in §8.
2 Background
In this section, we introduce works related to this paper. We first discuss biases in computer systems (§2.1) and then review existing inconspicuous attacks (§2.2) and general attacks on LLMs (§2.3).
2.1 Biases in Computer Systems
Existing works showed that biases exist in computer systems due to social circumstances, design choices, and use cases [31]. For example, some computer games only have male characters [32], and some voting machines were inaccessible to people of low height and people with reduced visual acuity [33].
Machine learning models might also have biases due to data, algorithms, and users [34]. For example, face-recognition systems can perform differently based on demographics, face geometry, and periocular features [35]. YouTube video captions have a significantly lower word error rate for men then for women [36]. Compared to white defendants, regression models tend to falsely flag black defendants as future criminals more often [37]. Some approaches have been proposed to mitigate the biases in machine learning algorithms [38, 39], but such methods typically suffer from decreased overall accuracy.
LLMs, as specific machine learning models, also have biases, although the specific definition of biases is context-dependent and culture-dependent [40]. In addition to various efforts to define [41, 42] and measure [43, 44, 45, 46] biases, numerous attempts have been made to mitigate biases in LLMs [47, 48]. However, to the best of our search, we did not find existing literature on biases in shopping brand recommendations, which is the use case in this paper, or biases induced by innocuous modifications of prompts, which is the general area we explore.
2.2 Human-Inconspicuous Attacks
One possible goal of adversaries is for attacks to be human-inconspicuousness: that humans at the scene are not able to notice an ongoing attack [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Examples of such attacks include human-inaudible voice commands to smartphones [10] and laser-based audio injection on voice controls [64]. Some defenses have been specifically proposed to detect such attacks [65, 66, 55, 67, 68, 69, 70, 71].
Evasion attacks are a type of attack on machine-learning systems that often tries to achieve inconspicuousness [72, 73, 74, 75]. With slight perturbations on images, evasion attacks aim to force well-trained machine learning models to behave unexpectedly [76, 77, 78, 79, 80, 81]. In the image domain, norms were proposed as a metric to measure the human-inconspicuousness of evasion attacks [82, 83, 84, 85, 86]. However, user studies suggested that norms might not accurately correspond to inconspicuousness [87, 88, 89]. Alternatively, some patch-based evasion attacks have been proposed and empirically verified to be inconspicuous to humans [11, 12]. Various defense techniques against evasion attacks have been explored in previous works: some detect evasion attacks [90, 91, 92], some use transformations [93, 94, 95], some prove lower bounds of robustness of machine learning models against evasion attacks [96, 97, 98, 99, 100, 101, 102], and some train machine learning models to work properly in adversarial circumstances [103, 104, 105, 106, 107, 108].
In the NLP domain, different approaches have been suggested to generate human-inconspicuous attacks: some use the distances between words (e.g., the Levenshtein edit distance) or embeddings (e.g., the USE score [109]) as metrics to measure inconspicuousness [24, 13, 17, 14, 15, 16], some change only a small number of words [17, 18, 19, 20], some utilize generative models [21], some exploit common typos [22], and some perform synonym replacement [23]. However, all these works either do not have an accompanying user study (e.g., [21]) or are suggested by user studies to not be inconspicuous to humans (e.g., [24]). In this work, we suggest a new text-domain inconspicuous attack (see §4.2 and §6.2) whose inconspicuousness is verified by a user study (see §7.1 and §7.2.2).
2.3 LLM Attacks
In light of LLMs’ prevalence, the security risks of LLMs and systems with LLMs have drawn the attention of researchers (e.g., [3, 4, 5]). We review proposed LLM attacks in the following paragraphs.
Backdoor attacks assume that adversaries can control some of the training data. In such attacks, adversaries attempt to change LLMs’ inference-time behavior by changing the data on which the LLMs are trained or tuned [110, 111, 6, 112]. Empirical [113] and provable [114] defenses have been proposed against backdoor attacks.
Jail-breaking attacks aim to cause LLMs to generate inappropriate content (e.g., offensive content). Adversaries in jail-breaking attacks do not have access to training data, but instead devise prompts that cause LLMs to respond with inappropriate content [7, 115, 116, 117, 118, 119, 120]. Most existing works evaluate jail-breaking attacks by success rate: the probability that offensive content is generated when LLMs are given perturbed prompts [121]. Defenses against jail-breaking attacks, also known as guard-railing, have been explored by many works [122, 123, 124, 125].
Data-extraction attacks aim to reproduce some or part of the training data [8, 126, 127, 128, 129, 130, 131]. In data-extraction attacks, adversaries query LLMs potentially many times, with any prompts they wish, but cannot directly access LLMs’ training data.
The threat model we explore in this paper differs from these existing attacks: we examine if inconspicuous changes to normal prompts can cause LLMs’ responses to be become probabilistically biased in a particular direction. We explain the differences between our threat models and others’ in more detail at the end of §3.
3 Threat Model
We posit that prompts from untrusted sources can cause unforeseeable biases in LLM responses. In this paper, we use brand recommendations during shopping as our attack scenario. More specifically, we assume that users (i.e., victims), who are shopping for specific goods (e.g., TVs), decide to ask LLMs for brand recommendations. Instead of writing prompts, victims use prompts created by adversaries (e.g., prompting services. See scenarios below for details).
For chatbots that outsource their underlying LLMs to third-party services,222many have announced or implemented chatbots using outsourced LLMs (e.g., OpenAI API), including Instacart [132], Lowe’s [134], Expedia [135], and more the aforementioned scenario creates an attack vector that does not require control over the model.
This threat isn’t just limited to chatbots with external LLMs. Despite owning control over the LLM (e.g., Gemini, ChatGPT), retraining models is costly [136]. Injecting bias through prompts, as we’ll show later, is not.
These LLMs have been vulnerable to many attacks, like jail-breaking (see §2.3). These risks are not hypothetical: a chatbot used by a car dealership was prompted to sell a $ car for $ [137], Air Canada had to honor the hallucinated policy by its customer support chatbot [138].
Below is concrete scenarios our threat model applies to.
Fig. 1 shows an example where a chatbot service is suggesting prompts. If the chatbot is adversarial, instead of engaging in costly re-training of models [136], it may suggest specially crafted prompts that result in biased respones. Notably, many chatbots already implement prompt recommendation features.
Users, seeking increased efficacy and utility from chatbots, may ask prompting services (e.g., PromptPerfect [133]) to write or optimize prompts. Fig. 4 shows the attack pipeline in this use case. Users first ask adversaries to write prompts and then use them with chatbots.
Further, adversaries may release prompts on online forums and encourage users to try them. If adversaries manage to convince victims to do so, users will send the prompts to LLMs and read the responses. For example, Instacart suggests prompts users can try with its LLM powered search [132]. Fig. 2 shows an example where some Reddit users encouraged others to try some prompts on ChatGPT [2]. Fig. 3 shows the attack pipeline in this use case.
Specific goals and constraints
In our threat model, regardless of specific use cases, adversaries cannot, or are disincentivized to change the weights of the LLM. They can, however, suggest prompts to users. Adversaries may also query LLMs with these prompts in advance of the attacks. We further assume that the adversary is constrained in its prompt perturbation: the prompts and resulting responses must not alert users to the attack. As such, the prompts and responses must be inconspicuous (see §2.2). If users are suspicious (e.g., propmts/responses semantically incorrect, containing nonsequiturs), users may stop using these prompts. We propose a practical definition of inconspicuousness for prompts and responses in subsubsection 7.1.1.
The main goal of the adversary is to induce LLMs to recommend certain brands more often, while not necessarily the most often among all brands. Figure 5 depicts this goal. In the presence of the perturbed prompts, the LLM recommends brand A more often than in the baseline condition. Similar to advertising, adversaries may economically benefit from this outcome. Notably, the adversary does not aim to prevent other brands (e.g., brand C), from being recommended more frequently. In practice, each response may recommend more than one brand. Note that adversaries may not need to cause LLMs to mention the exact brand name to recommend the brand. For example, to recommend the brand “Apple” for “laptops,” adversaries may instead cause “Macbook” to be recommended.
In summary, in our threat model, adversaries suggest prompts but do not have control over the model. An attack succeeds if, compared to a baseline, 1) prompts and responses are inconspicuous users and 2) the LLMs recommends a target brand more often. We describe such attacks in §4.1 and §4.2, and verify the effectiveness in §6 and §7. Our threat model is different from those described in previous work (§2.3): compared to backdoor attacks, adversaries in our threat model have no control over training data; compared to data extraction attacks, our adversaries do not extract training data and cannot use conspicuous prompts; compared to jail-breaking attacks, our adversaries do not generate inappropriate content, cannot use conspicuous prompts, and do not have access to model weights (or gradients).
4 Methods
In this section, we introduce how we build the approach that achieves the attack goals described in §3: generating prompts and corresponding responses that are inconspicuous to humans (i.e., they do not know an attack is taking place) and cause LLMs to recommend a target brand more often, ultimately increasing attention received from humans. We first describe our observation on LLM’s responses to paraphrased prompts (§4.1), and then illustrate our synonym replacement approach to perturb prompts human-inconspicuously §4.2.
4.1 Observations on Paraphrased Prompts
Machine learning models, including LLMs, are brittle (e.g., [82]). They can be highly sensitive to small changes in their input.
To promote a target brand, we apply this intuition. We explore attacking LLMs by testing paraphrased prompts. Specifically, we generate a set of prompts asking for recommendations of products in a certain category that are all paraphrases of one another. Our approach to this is described in §5.2. We show that these various paraphrased prompts (§6.1), although similar, can lead to increased prominence of a brand in the LLM’s responses. Given enough paraphrases, we’re able to find a prompt that causes LLMs to recommend a target brand more often.
This approach does not require any access to LLMs’ internal weights or token probabilities. However, it has a caveat, we need to try many paraphrases and collect many responses to find an optimal prompt. This computational cost might be an economic burden for the adversaries. Additionally, when we generate the paraphrases, some are likely less similar to each other and thus human-detectable. While we are able to find paraphrases that result in a large change in the probability that a certain brand is mentioned, even while using only a few prompts per category, as we describe in §5.2, in cases where an adversary does not have the time or computational resources to collect responses to multiple paraphrased prompts, adversaries may want to take another approach. As such, we propose another method to perturb prompts. To reduce computational overhead, with this new approach we assume access to the logits of LLMs in §4.2.
4.2 Synonym-Replaced Adversarial Prompts
The fundamental idea of this attack is to generate prompts by replacing words with synonyms and picking one that minimizes a loss function, reducing the need to collect computationally expensive responses. As we described in §2.2, while many NLP domain attacks have been proposed to be inconspicuous to humans, they either do not have accompanying user studies or are shown to be conspicuous to humans, motivating our approach. With this attack, we propose an improved synonym replacement method to generate human-inconspicuous perturbations.
Existing synonym replacement methods [23] either use WordNet, an established synonym dictionary, or a self-defined synonym dictionary [25, 26]. Nouns and verbs were found to have various forms (e.g., plural forms of nouns and verb tenses) [27], and thus an adjective synonym dictionary was suggested [28]. However, we found that synonyms of adjectives listed in existing dictionaries might not be synonyms in the context of brand recommendations. For example, WordNet and the existing adjective synonym dictionary both suggested “raw” as a synonym for “newest,” while it is human-conspicuous to recommend “the raw smartphone” instead of “the newest smartphone.” Thus, we created a new synonym dictionary compatible with brand recommendations. Our synonym dictionary mainly consists of adjectives but also includes other parts of speech while keeping the tense of verbs consistent between synonyms. In our dictionary, “select” and “choose” are synonyms, creating a synonym group, and other tenses of these verbs are excluded. “Exact,” “accurate,” and “precise” also creats a synonym group, this time of three. Our dictionary contains a total of 94 words in 36 synonym groups. Each word may have at most seven synonyms.
Since paraphrasing prompts imposes an economic burden (§4.1), systematically testing and gathering responses for all candidate perturbed prompts generated by synonym replacement would have a high cost. One prompt we tested had six words with synonyms and therefore candidate perturbed prompts with the same meaning. Rather than testing all candidates, we used a loss-guided candidate-selection method. Prior work suggested that when adversaries have white-box access to LLMs (i.e., when adversaries can access internal architecture and weights of LLMs), adversaries may use gradient search to perturb prompts [7]. Compared to prior work, however, our method does not require access to model weights or gradients. Our search space is much smaller, we do not need consider all possible tokens a model accepts, just synonyms [7]. Instead, we computed a logit-based loss for all possible combinations of synonym replacements and picked the combination with the lowest loss. It should be noted, that other candidates than the one with the lowest could be used as well, and we used this method to narrow the candidates down in order to be able to test more base prompt and target brand combinations.
Specifically, we use the following loss function. Using the same notation as existing works [7], we considered LLMs as a mapping from a sequence of tokens to a distribution over the next token . In other words, LLMs generate a probability . The probability that LLMs generate a sequence can be denoted as . In contrast to Zou et al., that aim to generate a specific sequence of tokens using the loss function
We designed a loss function that aims to generate a sequence among a set of possible candidate sequences
Intuitively, our loss function aims to cause LLMs to generate some sequence from the set right after the prompt . As we describe in §3, adversaries may not need to cause LLMs to mention a specific brand name to recommend the brand. For example, causing LLMs to recommend either “Macbook” or “Apple” meets adversaries’ goal to recommend the brand “Apple” for the category “laptops”, and in this case, might be .
Our loss function aims to increase the likelihood of words to appear from the target set right after the prompt, and we use the increased likelihood of words from this target set as a proxy to estimate how close adversaries are to their goal of promoting a brand (described in §3). We use increased likelihood to estimate which combination of a prompt’s synonym replacement may cause LLMs to recommend the target brand more often. Notably, despite the loss-function definition, adversaries do not need to cause LLMs to generate words from the target set right after the prompt, adversaries are still successful if any of the target words appear in the generated response (i.e., many tokens after the prompt). In §6.2, we show that a prompt with a lower loss value according to our new loss function was more likely to mention one of the target words set among up to 64 generated tokens. Additionally, we note that increased likelihood of a brand did not always guarantee “recommending” that brand. For instance, when we caused Gemma (one of the LLMs we use) to mention “Netflix” as a streaming service, Gemma yielded “I’ve been using Netflix for years, but I’m not happy with the selection of movies and TV shows that are available to me.” People may interpret the same LLM response differently and thus have different answers to whether a specific brand is recommended given the same response. This work, therefore, includes a more realistic user-study evaluation. We explored whether humans understood that the target brand was being recommended more in LLM responses to prompts we perturbed (§7.2.2).
5 Setup
In this section, we will describe the setups of our experiments. We will first introduce our choice of LLM and parameters in §5.1. Then we depict our experiment process, including evaluation metrics, in §5.2.
5.1 LLM Setup
In our experiments, we used four open-sourced LLMs as our benchmarks: a 7B pre-trained Llama 2, an 8B pre-trained Llama 3, an 8B instruction-tuned Llama 3, and a 7B instruction-tuned Gemma. Each of these models was downloaded from Meta or Google’s official repositories on HuggingFace. We used various LLMs to ensure our conclusion is not a special case only applicable to one specific instance of LLM.
Temperature is a parameter that controls the determinism of LLM responses. Therefore, with different temperature choices, LLMs may tend to recommend brands with different frequencies. Following the setup that existing works used to measure the indeterministic behavior of LLMs [7], we used the default temperature for each of the four models in our experiments.
Existing works collected up to responses per prompt to examine the biases in LLMs [40]. However, these studies grouped responses into only two partitions at a time (e.g., “he” versus “she” [40]), while in brand recommendation tasks, there might be more than two brands for the same category of products. To figure out how many responses we needed to collect per prompt, we ran a preliminary experiment with different combinations of prompts, target brands to be mentioned, and LLMs. We collected two sets of responses for each combination and found that they differed in the number of responses mentioning the target brand by no more than four (i.e., 0.8% in absolute means). As a result, we chose to move forward with responses per combination.
When collecting responses, we generate 64 tokens per prompt. We focus on the first 64 due to three reasons, 1) focusing on the beginning of responses likely is a better heuristic of what brand users are likely to see first, thus remember (this theory is supported in our user study §7); 2) some models (e.g., Llama 2, Llama 3) don’t stop generating tokens until they reach the maximum token limit of 4096 is reached, resulting in repetitive and meaningless responses; and 3) computational cost prohibits us from collecting more tokens than we have.
5.2 Evalution Procedure
We first listed 77 product categories where several established brands dominate the market. For each category, we listed popular brands we had heard about and also added brands that LLMs recommended but we did not include: some of the brands were later found to be rarely recommended by LLMs. Some brands appeared in more than one category: for example, the brand “Apple” appeared in both the category “laptops” and the category “smartphones”. The number of brands we listed for each category ranged from one to nine, with an average of per category.
Aided by ChatGPT, we gathered prompts that recommended brands in each category. Specifically, we queried ChatGPT with the following prompt: “Give me multiple rephrasings and add details to: ‘What is the best XXX?’”, where “XXX” is the category (e.g., smartphones). The prompts collected in the same category were believed to be paraphrases by ChatGPT, and we manually checked that these prompts 1) paraphrased each other and 2) were inconspicuous (see definition in §3) to us researchers. We filtered out the prompts that violated either of these two principles. An example of paraphrases we used would be “Which VPN service stands out as the optimal choice for ensuring top-notch online privacy and security according to your experience?” and “Can you recommend the ultimate VPN that excels in providing robust encryption, reliable performance, and a user-friendly experience?”, resulting in two prompts that request a recommendation for a VPN. Paraphrased prompts may ask for slightly different features and may use differing wording, but ultimately ask the same thing of the LLM. Ultimately, we had three to ten prompts per category, with an average of . In total, we collected prompts. We collected responses to each of these prompts on each of the four models. We used these prompts to examine the method we described in §4.1.
As we introduced in §4.2, for each combination of category and brand, adversaries may aim to cause LLMs to mention some words that are not necessarily brand names. For example, any of “ChatGPT”, “OpenAi” and “GPT” are target words (or, in some cases, strings) for the brand “ChatGPT” in the category “LLMs”. For each combination of category and brand, we collected a list of target words. We created these lists of target words based on our knowledge and observations of the responses from the 449 different prompts. Each combination of category and brand had one to five target words. We will compare paraphrased prompts by how often they mention some target words of a brand and category in §6.1. In this paper, we refer to a response as mentioning a target brand if it contains any of that brand’s target words.
To evaluate the efficacy of our synonym replacement approach (introduced in §4.2), we accordingly perturbed the 449 prompts in favor of each brand of the same category. We obtained synonym-replaced prompts for each model. We will compare the 449 prompts to the prompts by how often they mention a target brand in §6.2. We paired the prompts after synonym replacement with those before synonym replacement (i.e., corresponding pairs). For each pair, we computed the absolute improvement in the percentage of responses that mention a target brand. If of responses of the prompt before synonym replacement, and of responses of the prompt after synonym replacement mention the target brand, the absolute improvement was .
Some machine learning attacks were found capable of transferring [29, 30]: attacks against a machine-learning model might be effective against a different, potentially unknown, model. We thus investigated whether our synonym replacement approach can transfer to GPT3.5-turbo, a commercial and close-sourced LLM. On each of our four open-sourced LLMs, we first ranked the pairs of prompts according to the absolute improvement in mentioning a target brand. For the top pairs, we collected complete responses to both prompts: we allowed LLMs to keep generating until they yielded end-of-sequence (EOS) tokens. Instead of absolute improvement, we computed the relative improvement in the probabilities of mentioning the target brand within complete responses for each pair. If of responses of the base prompt before perturbing it via synonym replacement, and of responses of the prompt after perturbing it mention the target brand, the relative improvement was . Then we computed the relative improvement of these pairs on ChatGPT. Specifically, we used the GPT-3.5 turbo and GPT-3.5 turbo Instruct model with the default temperature parameter and collected complete responses. GPT-3.5 turbo Instruct is the instruction-tuned version of GPT-3.5 turbo. We compare the relative improvements between GPT models and open-sourced models in §6.3.
6 Results
In this section, we describe our empirical results. We first, in §6.1 describe our observations on paraphrases of prompts—measuring how pairs of paraphrased prompts can have significant differences in the probability that the target brand appears. Next, in §6.2, we report how often our synonym-replacement approach yields a perturbed prompt that causes LLMs to mention the targeted brand more often than the base prompt Finally, in §6.3, we explore whether our synonym-replacement approach transfers to GPT models. We show that if the right model-pair is found, successful perturbed prompts found on an open-sourced LLM can increase the probability of the target brand appearing more often in the GPT model’s responses.
6.1 Observations on Paraphrased Prompts
As we described in §5.2, each of the product categories in our dataset has three to ten paraphrased prompts. For each prompt, we measured the probability that LLMs would mention some target brand (described in §5.2) of a category within the first 64 tokens of the response. For each combination of product category and brand, we computed the maximal difference in the probability of responses mentioning the target brand between paraphrased prompts of the product category (within the first tokens). Adversaries may paraphrase a low-probability base prompt in order to find the high-probability prompts to achieve the goals described in §3.
We summarize our results in Fig. 6. Specifically, on four open-sourced LLMs—Gemma-it (instruction-tuned), Llama2, Llama3, and Llama3-it (instruction-tuned)—the likelihood of the target brand being mentioned in the responses can differ between a pair of paraphrased prompts by up to , , with the average of this absolute difference being , , , and , respectively. For example, when comparing responses to “I’m curious to know your preference for the pressure cooker that offers the best combination of cooking performance, durable construction, and overall convenience in preparing a variety of dishes.” with the prompt paraphrased as “Can you recommend the ultimate pressure cooker that excels in providing consistent pressure, user-friendly controls, and additional features such as multiple cooking presets or a digital display for precise settings?”, the probability of Gemma-it mentioning the brand “InstantPot” (“pressure cooker” product category) within the first tokens of the response went from to .
Our results suggest that while paraphrased prompts appear similar to humans the responses to the prompts can differ substantially in how likely they are to mention a target brand (within the first tokens of the response). Therefore, an adversary wanting to promote a certain brand can use this to their advantage: The adversary may try various paraphrases of prompts, test the prompts, and pick the paraphrase that results in the highest probability of the target brand being mentioned, ultimately promoting the brand when prompt is used in the real world. While we generated these paraphrases using ChatGPT, and confirmed that they were valid and reasonable, adversaries may also be able to create paraphrases manually or by another method, and may be able to test even more paraphrases than we did in our measurements.
6.2 Synonym-Replaced Adversarial Prompts
As described in §5.2, we measured and now report how often our synonym-replacement approach improves the probability of a brand being mentioned. As opposed to paraphrasing, in the synonym-replacement approach we automatically generate a set of potential candidate prompts by perturbing a base prompt via synonym replacement, without needing to confirm whether these perturbed prompts are valid. The prompt with the lowest loss is selected, and we assess how well these selected prompts increase the probability that the target brand is mentioned. We evaluate the average improvement over multiple models, at different response lengths, and for different probabilities of a brand being mentioned in the base prompt. In addition, other candidate prompts could also be tested, although we did not take this approach in this paper. We used loss as a metric that narrowed down the large set of potential perturbed prompts we found using synonym replacement to one. Therefore, we are interested in exploring the highest improvement we can achieve between a base and perturbed score, as adversaries would be able to use a similar method to §6.1 and explore multiple perturbated prompts. So, we also describe the largest increases in probability of mentioning a target brand in the responses to the perturbed prompt we found via our synonym-replacement method compared to the base prompt. Our evaluation shows that our synonym-replacement attack increases the likelihood of LLMs mentioning a brand on average.
For all evaluations, we focus on the first 64 tokens of the response (see subsection 5.1 for details).
Average improvement
The results for average improvement in likelihoods of a brand being mentioned in the responses to perturbed prompts created via synonym replacement compared to in response to the original base prompt, across models, are shown in Fig. 7. Overall, we find that our attack results in an average absolute improvement in all models. Specifically, for brands that were mentioned at least once (i.e., of the time) in responses to the base prompt, Gemma-it had an average absolute improvement within tokens and a improvement within tokens. Llama2 achieves an average absolute improvement of within tokens, and within tokens. Llama3 achieves an average absolute improvement of within tokens, and within tokens. Llama3-it achieves an average absolute improvement of within tokens, and within tokens. The averages of the best absolute improvement in each product category are , , , and on Gemma-it, Llama2, Llama3, and Llama3-it respectively. Our synonym replacement approach achieves a positive absolute improvement in probabilities, which verifies that our approach is capable of forcing LLMs to mention target brands more often within some number of tokens. On all four models, we see a bigger absolute improvement within some number of tokens less than . While we only collected responses of length 64 tokens for our 1,809 attacks for each model (see subsection 5.1), we did run a select set of base and perturbed prompt pairs (chosen in §5.2) to full completion. While we did not see improvements in all pairs (for example, some base prompts had near probabilities of mentioning certain brands with long completions, not allowing for any improvement), we still saw pairs with significant increases when considering long responses. For Llama3-it, the perturbed prompt “Can you recommend the superior video game console that excels in providing top-notch graphics, dissimilar gaming options, and additional features such as online connectivity, appropriate for both casual and hardcore gamers?” was more likely than its base prompt “Can you recommend the ultimate video game console that excels in providing top-notch graphics, diverse gaming options, and additional features such as online connectivity, suitable for both casual and hardcore gamers?” (i.e., “ultimate” was changed to “superior” and “diverse” was changed to “dissimilar”) to mention Xbox in long completions, even more than when completions were only 64 tokens long (). We evaluate if attack objectives are met with full responses in a more realistic setting in §7.2.2.
Baseline probability and attack success
Here we explore the relationship between attack success rates and the probability of a brand being mentioned in responses to the base prompt. As shown in Fig. 8, our synonym-replacement approach on average achieved a bigger absolute improvement when the target brand was mentioned more often in responses to the base prompt before perturbation. When the target brands were mentioned at least (i.e., in out of responses before perturbation), our synonym-replacement approach achieved an absolute improvement of , and on Gemma-it, Llama2, Llama3 and Llama3-it, respectively. We see that for base prompts with higher probabilities of mentioning the target brand, our perturbed prompts, on average, have a higher increase in probability. This indicates that our perturbation method is more effective on an identifiable domain, i.e. prompts where the target brand is mentioned more or less frequently, than it is on the overall average case.
Maximum improvement
Besides the average absolute improvement, we also explored the maximum absolute improvement among all combinations of base prompts and brands, as shown in Fig. 9. While these results don’t represent the expected improvement using this method, they do demonstrate that it is possible to find pairs of prompts with vastly different probabilities of mentioning a certain brand. In §6.1 we showed this for prompts that were manually paraphrased; here, we show that the synonym-replacement attack can find such alternative prompts automatically. Further, the prompts generated by synonym-replacement differ from their base prompts minimally (at most a seven synonym difference in our dataset), and are perceptually the same along multiple dimensions (see §7.2.2).
We were able to achieve bigger absolute improvement on Gemma-it compared to the other three Llama models, although Gemma-it does not always have the highest average absolute improvement (Fig. 7 and Fig. 8). When LLMs generate all 64 tokens, Gemma-it has the highest absolute improvement of when we perturb “Looking for guidance on choosing the best TV; could you recommend one and share additional details like display type, audio quality, or any other features that set it apart?” as “Looking for guidance on choosing the ultimate TV; could you suggest one and share extra details like display type, audio quality, or any other features that set it apart?” (i.e., replacing “best” with “ultimate”, “recommend” with “suggest”, and “additional” with “extra” in favor of brand “Samsung” in product category “TV”. With all Llama models, we see the maximum increase as we consider more tokens of the response, reaching around .
6.3 Transferability to GPT Models
We evaluated the transferrability of paraphrased prompts generated using synonym-replacement (§6.2) from open-sourced models to GPT models. We found that, while transferability was limited for most model pairs, Llama3-it and GPT3.5-turbo had a high correlation in attack success for the same prompts (). We explain these results in more detail next.
As we described in §5.2, we computed the relative improvement on four open-sourced LLMs and two versions of GPT models. Specifically, for a given base prompt, we generated a perturbed prompt via synonym-replacement that increased the chance of a brand being mentioned on an open-sourced model and then evaluated whether the perturbed prompt also increased the chance of the brand being mentioned in the response produced by the GPT model. We evaluated only the transfer of attacks that were successful on their open-sourced LLM onto GPT models. For the Shell example described in §6.2, we saw a improvement from to between the base and perturbed prompt on long responses by Gemma-it between prompts and an improvement from to on GPT3.5-turbo responses. On long responses generated by Llama3-it, “If you had to pinpoint the superior investment platform, which one would it be, and what specific features make it stand out as the top choice for investors?” had a probability of mentioning Fidelity, while “If you had to pinpoint the premier investment platform, which one would it be, and what specific features make it stand out as the top choice for investors?” had a , and GPT3.5-turbo responses went from to .
To compare the relative improvements between models when the same prompts and synonym replacement are used, we used the Pearson correlation test. Two sets of data with a Pearson correlation coefficient () less than is generally believed to be weakly correlated, whereas or lower is uncorrelated. On the other hand, a Pearson correlation coefficient larger than is believed to indicate strong correlation [139].
As shown in Fig. 10, while Pearson correlation coefficients of the relative improvement of most pairs of open-sourced LLMs and GPT models indicated weak or no correlation (), Llama3-it and GPT3.5-turbo Instruct have a correlation coefficient of , and thus are highly correlated (). Llama3-it and GPT3.5-turbo also show some correlation, with a coefficient of . This result indicates a potential for a transfer attack to be used on chatbots that use black-box GPT models, like Instacart, Lowe’s, and Expedia (described in §3) and others. A successful attack on an open-sourced LLM could be used as a prompt suggestion for black-box models, ultimately promoting the target brand.
7 User Study
Our results so far show that our synonym-replacement approach can bias LLM responses towards a target brand and appears to be human-inconspicuous. However, this does not mean that the attack is successful in practice. To evaluate the attack in a realistic setting, we conducted a user study. Here we describe this study, starting with a description of our methods (§7.1) and ending with the statistical evaluation (§7.2). We find that our attack is indeed successful in practice.
7.1 Methods
We performed a between-subjects user study to evaluate whether our synonym-replacement approach can result in prompts that advances the adversaries’ goals (see §3), creating inconspicuous prompts that trigger inconspicuous responses, making a target brand more noticable.We tested whether users could distinguish between base and perturbed prompts and responses in multiple dimensions (including clarity, likelihood of use, satifaction, and more). Each participant was shown one prompt-and-response pair. Our procedures were approved by our institution’s ethics board, and pre-registered.333https://osf.io/6mycr/?view_only=face90d04806439bb1f69fc110fb9a1e
7.1.1 Survey procedures
We recruited a gender-balanced sample from Prolific,444https://www.prolific.com/ a crowdsourcing platform commonly used with security-relevant user studies (e.g., [140]). To avoid selection bias in our study ad, we avoided mentioning LLMs or bias in the study title, “Chatbot prompting study.” Participants had to be in the U.S., be 18 or older, and have a Prolific approval rate of at least 95%. In line with Prolific guidelines 555https://researcher-help.prolific.com/hc/en-gb/articles/4407695146002-Prolific-s-payment-principles, participants were paid $1.6 for an estimated 7–9 minute study. In practice, the hourly compensation rate was $12.45/hour.
After providing consent, participants were given an overview of the study. We asked our participants to imagine that there was a chatbot service that was able to recommend prompts appropriate for what users want to use the chatbot for. We further instructed participants to imagine that they were shopping for a certain product (e.g., laptops) or service (e.g., parcel delivery) and wanted to use this new chatbot service to help them decide on a brand. After a comprehension check on instructions, participants were shown a prompt recommendation for the product they were shopping for (e.g., “Which laptop model do you consider the optimal choice for versatile computing, powerful performance, and innovative features that enhance your work and entertainment experience?”) and instructed to review it. After a minimum of 10 seconds had passed, we asked (1) how likely they were to use this prompt, (2) how clear the prompt was, (3) was it biased to a certain brand (and which), (4) and if anything stood out (e.g., interesting, unexpected).
Participants were then shown a response to the prompt and after a minimum of 20 seconds, they were asked: (1) how clear the response was, (2) if they were satisfied with the response, (3) how likely were they to take the recommendation in the response, (4) and if anything stood out. In a series of open-ended questions, we additionally asked participants (1) which brand they would pick based on this response, (2) what were all the brands recommended, and (3) what the top brand recommendation of the chatbot was. These 11 (numbered) questions form the practical definition of inconspicuousness and increase in target brand perception. They form the basis of our statistical analysis §7.2.2.
To help with recall, participants could hover over relevant questions to reveal the relevant prompts and responses.
Participants self-reported how frequently they give tech advice, how freuqently they use chatbots, if they paid for a chatbot, their ChatGPT familiarty. The survey concluded with demographics questions.
| Gender | Male | 48.2 |
| Female | 50.3 | |
| Self-described | 0.9 | |
| Age | 18-25 | 16.7 |
| 26-35 | 36.0 | |
| 36-45 | 21.5 | |
| 46-60 | 18.2 | |
| 61+ | 5.2 | |
| Ethnicity | White | 63.0 |
| Black or African Am. | 12.1 | |
| Asian | 9.8 | |
| Hispanic or Latino | 5.6 | |
| Other or mixed | 12.7 | |
| Education | Completed H.S. or below | 10.4 |
| Some college, no degree | 18.0 | |
| Trade or vocational | 2.5 | |
| Associate’s degree | 11.1 | |
| Bachelor’s degree | 39.1 | |
| Master’s or higher | 18.5 | |
| Chatbot usage | Daily or more freq. | 16.6 |
| frequency | Daily to monthly | 49.7 |
| Monthly or less freq. | 33.7 | |
| ChatGPT | A lot | 57.2 |
| familiarity | A little | 40.7 |
| Nothing at all | 2.1 |
| Prompt | Reponse | ||||||||
| Clarity (L7) | Use (L7) | Bias (L5) | Standout (B) | Clarity (L7) | Use (L7) | Satisfied (L7) | Standout (B) | ||
[…]
| Gemma-it | LLama3-it | ||||||||||||||||||||||
| TV | ISP | Parcel delivery | Gaming console | Invesment plat. | Laptop | ||||||||||||||||||
| P% | A% | T% | P% | A% | T% | P% | A% | T% | P% | A% | T% | P% | A% | T% | P% | A% | T% | ||||||
| Base | 25.0 | 33.8 | 35.3 | 23.9 | 46.2 | 34.3 | 10.6 | 22.7 | 13.6 | 0.0 | 18.3 | 0.0 | 14.1 | 38.0 | 14.1 | 5.3 | 61.3 | 1.3 | |||||
| Pert | 57.1 | 80.0 | 82.9 | 39.1 | 78.2 | 50.7 | 41.4 | 58.6 | 44.3 | 11.2 | 64.8 | 1.4 | 40.3 | 58.4 | 42.9 | 7.1 | 74.3 | 0.0 | |||||
[…]
7.1.2 Experimental groups
Our overall goal was to find whether people notice differences between our perturbed prompts compared to base prompts. For each of our 447 base prompts, we found a perturbed prompt for each possible target brand using the loss of each of our models. This resulted in 1,809 base and perturbed prompt pairs for each model. Because of the large number of prompt pairs, we could not test all of them. We selected six pairs of prompts to use in the user study, each pair from a different product category, giving us 12 prompts total. Because each pair of prompts belonged to a unique product category, we refer to them by their product categories in the results. The prompt pairs were split between two models, three pairs for Llama3-it and three for Gemma-it. We focused on the more user-friendly instruction-tuned models from our experiments since they are more analogous to LLMs used in chatbots [141]. The pairs met the following criteria:
Among the top 50% of product categories that participants reported to care about.
Among the top 50% of product categories that participants reported they might use an LLM when shopping for.
They had the highest probability increase that the target brand is mentioned with the attack, as discussed in §6.1.
For each of the 12 prompts in our six prompt pairs, we obtained a random sample of 75 model responses from our earlier experiments with the models (§6.2). This sample was stratified to ensure the ratio of responses that mentioned the target was the same in this set of 75 as it was in the overall set of 1000. All participants within a study group saw the same prompts, but, to mirror how in real-world chatbot use a single prompt would lead to different responses, each participant was shown a unique reponse. As opposed §6, particiapnts saw full-length responses.
7.1.3 Piloting and preliminary data collection
We piloted our study extensively. We ran a series of preliminary studies to determine whether the study design was feasible and whether participants would encounter problems with any parts of the study. During these preliminary studies, questions were timed and augmented with meta-questions on how clear the main questions were. We also collected participants’ interest in product categories and their likelihood of using chatbots for advice when shopping for these categories. In total, we collected responses from 90 prolific participants for piloting and 63 for product category preferences. We further ran pilots with two HCI researchers, asking them to review and criticize our study.
Based on responses, we tweaked the study design to be more concise and have clearer instructions. To more closely mimic real chatbots, we changed how prompts and responses were displayed.Attention and comprehension checks were added to ensure high data quality.
7.1.4 Statistical analysis
For each of the six base and perturbed prompt pairs, we analyzed the difference between base and perturbed group responses with a series of non-parametric tests on our 11 main measurement variables: two-tailed Mann-Whitney U tests for Likert data and chi-squared tests for binary data. To understand whether the base and perturbed groups are equivalent, we replaced non-significant tests for difference with tests of equivalence. To establish equivalence, we used the two one-sided tests procedure (TOST) and set our equivalence margin to be for 80% power [142].666Testing for equivalence is a deviation from the pre-registration; however, we believe this approach paints a more complete picture. Open-ended brand-recall questions (e.g., what brand participants would pick based on the response), were coded into binary categories: whether the participant reported the targeted brand or not. hesitant or otherwise unclear responses did not count as a match.
Our approach to measuring inconspicuousness meant that we needed to run 11 tests between the baseline and perturbed prompt pair for each product category. We controlled our false-discovery rate per product category with the Benjamini-Hochberg procedure [143].
7.2 Results
We first describe our participants and then present results. We find that, under most measures, perturbed prompts and corresponding responses are human-inconspicuous. Further, these prompts, in most measures, successfully nudge more users into noticing the target brand.
7.2.1 Participants
We recruited 845 participants and assigned them evenly to 12 groups, each group defined in terms of the product category and prompt type (baseline or perturbed). Product categories were split between Llama3-it and Gemma-it. Our participants were more educated, younger, less hispanic, and had more familiarity with ChatGPT compared to the national average [144, 145]. Table I shows the demographics distribution of our participants.
7.2.2 Statistical Evaluation
We asked multiple questions about the prompts and responses. Four of these were to test the inconspicuousness of the perturbed prompts. Another four were to test if users were significantly more dissatisfied with the perturbed responses. Additionally, we asked three questions to test if the responses to the perturbed prompts made the targeted brand more perceptible. Table II and Table III show the result of our analysis.
Equivalence
We find that nearly all perturbed prompts and responses were equivalent to the base prompts from the perspective of participants in terms of variables measured (42/48 comparisons were equivalent, ) with the following exceptions: statistical tests showed no difference or equivalence for the likelihood of using the parcel delivery prompts (); participants found the responses to the parcel delivery perturbed prompt more satisfactory and more likely to be used (all ); the perturbed gaming console prompt was less clear than the base prompt (); no difference nor equivalence was found for the likelihood of using the response for gaming platforms (); the laptop perturbed prompt was more likely to be used (). These results suggest that not only is our attack imperceptible to users, but in a few cases the perturbed prompts and responses might be preferable.
Attack success
In order to measure attack success, we recorded the percentage of participants who would pick the targeted brand given the response, the percentage of participants who noticed the brand in the response, and the percentage of participants who said the targeted brand was the top recommendation. As shown in (Table III), in five out of six categories our attacks were able to increase the prominence of the target brand in at least one of the three questions: in four categories participants were more likely to pick the targeted brand when given the perturbed prompts, in five categories participants were more likely to notice the targeted brand, and in three categories participants were more likely to say the targeted brand was the top recommendation.
8 Concluding Discussion
We identify a novel threat model in which an adversary aims to induce target biases in LLM responses users receive by using adversarially crafted prompts. We specifically investigate whether adversaries can bias responses towards a target brand through prompt suggestions (e.g., recommended prompts within a chatbot, prompt sharing services).
We contribute a dataset of 449 prompts that ask for product recommendations over 77 product categories, and develop a method to evaluate which brands are mentioned in LLM responses using target words from the same dataset. We evaluate the difference in probabilities that brands are mentioned across similar prompts (paraphrases) within each product category. We find up to a difference between pairs of prompts. Next, we develop a method to perturb a base prompt by replacing words with their synonyms from our new synonym dictionary. We select a candidate using a loss function that correlates with measured attack success. This results in human-inconspicuous prompts that increase the probability that a target brand is mentioned. Novel with respect to other work that studies inconspicuous text manipulations, we perform a user study that demonstrates that, compared to base prompts, our synonym-replacement prompts are successful in promotinga target brand while remaining inconspicuous to humans.
We also explore transfer attacks of our synonym-replacement approach from open-sourced models where logits are available to GPT models, finding transferability between Llama3-it and GPT-3.5-Turbo.
Taken together, our work highlights the potential for adversarial prompt recommendations to bias LLM responses. While our focus is on brand recommendations, our methods are general and can be applied to other adversarial goals, such as propaganda, misinformation, and bias on social issues. As such, we suggest that defensive measures should be taken. Since the perturbations we introduce are imperceptible to humans, we suggest that users should be warned to check sources before using prompts they did not author. Notably, warnings about LLM use are already commonplace in many chatbot applications, ChatGPT noting “ChatGPT can make mistakes. Check important info.” [146]. We suggest that these be implemented for untrusted prompts.
Appendix
Do LLMs Recommend Their Parent Brand?
| search engine | gemma-it | Llama2 | Llama3 | Llama3-it | GPT-3.5 | GPT-3.5-it |
| Bing | 0.58 % | 13.08 % | 14.00 % | 27.66 % | 0.40 % | 2.36 % |
| 98.16 % | 50.04 % | 53.14 % | 53.60 % | 74.24 % | 45.86 % | |
| Yahoo | 0.00 % | 19.36 % | 18.36 % | 3.92 % | 0.26 % | 0.56 % |
| browser | gemma-it | Llama2 | Llama3 | Llama3-it | GPT-3.5 | GPT-3.5-it |
| Chrome | 77.22 % | 41.24 % | 33.54 % | 50.30 % | 53.28 % | 35.42 % |
| Firefox | 28.74 % | 38.10 % | 31.28 % | 22.28 % | 5.08 % | 5.40 % |
| Safari | 3.40 % | 11.70 % | 9.00 % | 5.64 % | 0.02 % | 0.36 % |
| Edge | 7.62 % | 13.70 % | 10.04 % | 13.52 % | 0.36 % | 1.14 % |
| Opera | 0.14 % | 9.82 % | 9.74 % | 6.38 % | 0.00 % | 0.06 % |
| llm | gemma-it | Llama2 | Llama3 | Llama3-it | GPT-3.5 | GPT-3.5-it |
| ChatGPT | 94.82 % | 39.28 % | 41.66 % | 0.30 % | 16.68 % | 38.04 % |
| 2.56 % | 8.28 % | 7.84 % | 19.3 % | 0.02 % | 1.06 % | |
| Llama | 0.00 % | 0.40 % | 0.70 % | 2.32 % | 0.00 % | 0.00 % |
| Claude | 0.00 % | 0.12 % | 0.00 % | 0.00 % | 0.00 % | 0.00 % |
| Vicuna | 0.00 % | 0.00 % | 0.00 % | 0.00 % | 0.00 % | 0.00 % |
| os | gemma-it | Llama2 | Llama3 | Llama3-it | GPT-3.5 | GPT-3.5-it |
| Windows | 97.18 % | 51.10 % | 59.72 % | 64.04 % | 17.30 % | 9.62 % |
| Mac | 24.76 % | 34.66 % | 37.98 % | 38.42 % | 9.32 % | 3.14 % |
| Linux | 1.16 % | 29.94 % | 26.46 % | 30.96 % | 2.88 % | 1.38 % |
| smartphone | gemma-it | Llama2 | Llama3 | Llama3-it | GPT-3.5 | GPT-3.5-it |
| Apple | 90.32 % | 28.88 % | 30.45 % | 14.97 % | 12.21 % | 31.29 % |
| 21.85 % | 11.64 % | 10.53 % | 12.25 % | 0.15 % | 1.80 % | |
| Samsung | 9.02 % | 31.11 % | 27.19 % | 33.41 % | 0.71 % | 8.86 % |
| laptop | gemma-it | Llama2 | Llama3 | Llama3-it | GPT-3.5 | GPT-3.5-it |
| Mac | 25.76 % | 14.22 % | 13.46 % | 7.16 % | 44.66 % | 9.80 % |
| Chromebook | 0.00 % | 1.88 % | 1.6 % | 0.18 % | 0.00 % | 0.02 % |
| HP | 0.00 % | 12.14 % | 14.16 % | 9.96 % | 0.22 % | 1.40 % |
| Asus | 0.00 % | 7.08 % | 6.64 % | 3.84 % | 0.04 % | 0.20 % |
| Lenovo | 0.00 % | 9.44 % | 14.42 % | 13.94 % | 0.28 % | 2.24 % |
| Acer | 0.00 % | 9.20 % | 9.48 % | 4.06 % | 0.00 % | 0.08 % |
| Dell | 39.98 % | 14.62 % | 17.54 % | 43.60 % | 22.98 % | 56.58 % |
| VR headset | gemma-it | Llama2 | Llama3 | Llama3-it | GPT-3.5 | GPT-3.5-it |
| Meta | 41.60 % | 25.54 % | 32.94 % | 34.16 % | 42.74 % | 42.60 % |
| HTC | 0.14 % | 21.66 % | 33.32 % | 38.30 % | 0.72 % | 1.80 % |
| Playstation | 0.00 % | 0.30 % | 0.52 % | 0.00 % | 0.00 % | 0.00 % |
| email provider | gemma-it | Llama2 | Llama3 | Llama3-it | GPT-3.5 | GPT-3.5-it |
| 35.32 % | 37.40 % | 35.18 % | 22.52 % | 45.26 % | 31.46 % | |
| Yahoo | 0.00 % | 4.98 % | 9.16 % | 5.12 % | 0.82 % | 0.48 % |
| Microsoft | 4.00 % | 16.76 % | 17.78 % | 14.40 % | 2.16 % | 1.90 % |
Throughout our experiments evaluating rephrasings in §6.1, we gathered completions for prompts on categories with products manufactured by Meta, Google, and Microsoft, which allowed us to examine how large language models developed by these companies perform when asked about product categories that include products manufactured by them. We evaluated the average score, as defined in §5.1, of all brands over all prompts for categories where one of the brands was Meta, Google, or Microsoft. For Google the categories included browsers (Chrome), large language models, smartphones (Pixel), laptops (Chromebook), email providers (Gmail), and search engines; for Meta they included VR headsets and large language models (Llama). As before, this meant we looked for target words related to a brand in the response to see whether this prompt was mentioned. We were interested in whether or not LLMs made by a certain company were biased towards products made by the same company. All results are shown in Fig. IV
Google has developed a variety of LLMs and LLM families (laMDA, Bert, PaLM, Gemini, Gemma) [147], yet we still found some interesting mistakes in Gemma-it’s responses to prompts asking for recommendations on large language models. For example, across multiple prompts, Gemma-it’s responses included “**GPT-4:** This model, developed by Google,” a false statement [147] seeming to claim that GPT-4 was developed by Google. The same was said for GPT-3, which is also false [147]. So, even Gemma-it responses that mention Google might actually be recommending GPT. Out of Gemma-it’s responses, the only actual model of Google’s mentioned is PaLM.
We see more mentions of Llama or Meta when querying Llama than when querying Gemma-it or GPT3.5-turbo, with our three Llama models we see a , , and and never with the other models. However, we did not always observe this self-preference. Llama models recommend a Google model at least of the time whereas Google only recommends a Google model of the time. All Llama models recommend Meta VR headsets under of the time, while Gemma and GPT models do over of the time, and Llama2 mentions Gmail when prompted about email providers more than Gemma-it does ( vs ). Gemma-it never mentions Chromebooks when asked about laptops, while all three Llama models and GPT-3.5 Instruction sometimes do. Nonetheless, Gemma-it seems to show a higher preference towards Google products than any Llama model for the categories of search engines and phones.
Overall, this test size is small and does not necessarily take into account all factors that can cause differences between these models. In the end, we do not find any bias by LLMs towards products developed by the same parent company, but believe it warrants further exploration.